40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Aug 27 02:09:14 2015, ericq, Update, LSC, AUX X Laser Current Changed 
    Reply  Fri Sep 18 11:59:08 2015, ericq, Update, LSC, AUX X Laser Current Reverted 
Message ID: 11533     Entry time: Thu Aug 27 02:09:14 2015     Reply to this: 11619
Author: ericq 
Type: Update 
Category: LSC 
Subject: AUX X Laser Current Changed 

I spent some time tonight chasing down the cause of huge RIN in the X green PDH transmitted light, which I had started seeing on Monday. This was preventing robust locking, since the ALS sensing noise was ~10x worse above 50Hz, thus making the AO transition much flakier (though, impressively, not impossible!)

I went down to the X end, and found that turning the laser diode current down by 0.1A (from 2.0 to 1.9) smoothed things out completely. Unfortunately, this causes the power to drop, from GTRX of 0.45 to 0.3, but the ALSX sensitivity is unchanged, as compared with the recenent "out of loop" template. 

This also seems to have changed the temperatures of the good modes, as no beat was evident at the previously good temperature. Beats were found at +5400 and +10500 counts on the slow servo offset slider; I suspect the third lies around the edge of the DAC range which is why I couldn't uncover it. In any case, I've parked it at 10500 for now, and will continue locking; nailing it down more precisely and offloading the slider offset to the laser controller will happen during daytime work...

ELOG V3.1.3-